# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ IV-C Existence of the Codebooks and Achievability Bound
    $\begin{split}\log M_{n}=&nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))\\ +&\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}\\ \times&\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\\ +&O(\log n).\end{split}$ (98)
@@
    $\gamma_{1}=n(1+\rho)[\mathsf{C}(\mu,\Psi(n),\sigma^{2})+\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})]$ (100)
@@
    The power level $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$ with $c_{4}$ is any constant such that
    $\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})<\delta.$ (101)
Under the above conditions, we have
$\log M_{n}^{*}(n,\epsilon,\delta)\geq$ $\frac{\mu^{2}c_{4}}{2}\sqrt{n}\log e$
$+$ $\mu\sqrt{c_{4}}n^{\frac{1}{4}}\Phi^{-1}(1-\epsilon)\log e+o(n^{\frac{1}{4}})$ (102)
@@
$\log K_{n}=\begin{cases}\frac{1+\mu^{2}\sigma^{2}-2\mu\sigma^{2}}{2\sigma^{2}}\cdot c_{4}\sqrt{n}\log e+O(n^{1/4})\\ \,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<\max\{1,\frac{1}{2\mu-\mu^{2}}\},\\ O(1)\,\ \ \ \ \ \,\text{otherwise}.\end{cases}$ (103)
@@
By expurgating the $(1-\lambda_{2})M_{n}$ codewords with largest $\varepsilon_{sw}^{(2)}$ for all $s$ from each sub-codebook $\{\bm{x}_{sw},w\in\langle M_{n}\rangle\}$, we get a new family of codebooks, each of which has size $\lambda_{2}M_{n}$. As the average error probability of each old codebook is less than $\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)}$, from Markov’s inequality, the new family of codebooks should satisfy
$\underset{s,w}{\max}\varepsilon_{sw}^{(2)}\leq\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)}.$ (104)
For the condition of reliability, from Lemma 3, for sufficiently large $n$, the following inequality holds,
$\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{m,1}}\big{(}\frac{\bar{Q}^{(n)}_{m,1}}{Q^{(n)}_{m,1}}\big{)}\leq\frac{2}{n^{c_{0}-c_{1}-c_{2}}}<\epsilon-\epsilon_{1}.$ (105)
For the covertness condition in (96), from (101) and Lemma 3, the term $\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})$ satisfies
$(1+\iota^{\prime})$ $\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})$
@@
$\log$ $\big{(}1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}M_{n}K_{n}}\big{)}+An\cdot ae^{-bn}$
$+$ $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+\lambda_{3}+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})$
@@
$+$ $\frac{1}{n^{c_{3}}}+(1+\iota^{\prime})\mathbb{D}(\bar{Q}^{(n)}_{1}\|\bar{Q}^{(n)}_{0})$
$\leq$ $\delta$ (107)
@@
$\exp$ $\bigg{[}-M_{n}\cdot\bigg{(}\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})\bigg{)}\bigg{]}+\frac{1}{\lambda_{1}}$
@@
$\leq$ $1$ (112)
$\log M_{n}^{*}(n,\epsilon,\delta)\geq$ $\frac{n}{2}\log(1+\mu c_{4}\frac{\sqrt{n}}{n})+\frac{(\mu^{2}-\mu)c_{4}\frac{\sqrt{n}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})}\cdot\log e+\sqrt{n\cdot\frac{2\mu^{2}c_{4}\frac{\sqrt{n}}{n}+\frac{c_{4}^{2}\mu^{2}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})^{2}}}\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\log e$
@@
With $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get a lower bound of $\log M_{n}^{*}(n,\epsilon,\delta)$ in (IV-C). The condition in (99) is equivalent to
$\log M_{n}+\log K_{n}=\gamma_{1}+c_{5}\log n+1$ (116)
Consequently, the least number of bits of $K_{n}$ should satisfy
$\log K_{n}$
$=$ $\max\{0,\gamma_{1}+c_{5}\log n+1-\log M_{n}\}$
=$\overset{(a)}{=}$ $\max\{0,n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))$
$+$ $\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]$
-- $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}$
@@
=$\overset{(b)}{=}$ $\max\{0,n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))$
$+$ $\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]$
@@
$\begin{split}&n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))]\\ =&n[\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\frac{1}{2}\log(1+\mu\Psi(n))]\\ =&\frac{n}{2}\log\big{[}1+\frac{\mu\Psi(n)(1-\sigma^{2})}{\sigma^{2}(1+\mu\Psi(n))}\big{]}\end{split}$ (118)
Using $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$ , we get $\frac{1}{2}\mu\big{(}\frac{1}{\sigma^{2}}-1\big{)}c_{4}\sqrt{n}\log e+O(1)$. In addition,
$\begin{split}&n[\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]\\ =&\frac{(1-\mu)\Psi(n)(1+\mu\Psi(n)-\mu\sigma^{2}-\mu^{2}\Psi(n))}{2(1+\mu\Psi(n))(\sigma^{2}+\mu\Psi(n))}\log e\end{split}$ (119)
By using $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get the sum of (118) and (119) as
$\frac{(1-\mu)(1-\mu\sigma^{2})}{2\sigma^{2}}c_{4}\sqrt{n}\log e+O(1).$ (120)
and the term $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}=O(n^{1/4})$. Let $\phi(\mu,\sigma)=\frac{1+\mu^{2}\sigma^{2}-2\mu\sigma^{2}}{2\sigma^{2}}$, we discuss it in different cases.
  • If $\sigma^{2}\leq 1$, i.e., the channel between Alice and Willie is less noisy than the main channel, we have
    $\log K_{n}=\frac{1+\mu^{2}\sigma^{2}-2\mu\sigma^{2}}{2\sigma^{2}}c_{4}\sqrt{n}\log e+O(n^{1/4}).$ (121)
  • If $\sigma^{2}>1$, then $\phi(\mu,\sigma^{2})<0$ if $\sigma^{2}>\frac{1}{2\mu-\mu^{2}}$, i.e., the first order asymptotics of the optimal number of key bits should be $O(1)$ if $\sigma^{2}>\max\{1,\frac{1}{2\mu-\mu^{2}}\}$.
@@
$\log K_{n}=\begin{cases}\frac{1+\mu^{2}\sigma^{2}-2\mu\sigma^{2}}{2\sigma^{2}}\cdot c_{4}\sqrt{n}\log e+O(n^{1/4}),\\ \,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<\max\{1,\frac{1}{2\mu-\mu^{2}}\}\\ O(1)\,\ \ \ \ \ \,\text{otherwise}\end{cases}$ (122)
We can see that the size of the key depends on both $\mu$ and $\sigma^{2}$. If $\phi(\mu,\sigma)>\frac{2\mu\sigma-2\mu\sigma^{2}}{2\sigma^{2}}>0$ and $\frac{d\phi}{d\mu}=2(\mu-1)<0$. Since $\mu$ is the parameter that represents the truncation, this implies that smaller $\mu$ will result in larger size of key. In addition, $\phi$ is monotonically decreasing with $\sigma^{2}$, which is consistent with the following intuition. When the noise level at the channel between Alice and Willie is smaller, then Alice should use a larger size of key to confuse Willie.
@@
Note that, we choose $\Psi(n)=O(\frac{1}{\sqrt{n}})$ in the condition (d). In fact, any $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $\tau\geq\frac{1}{2}$ is qualified for the inequality (IV-C) if $n$ is sufficiently large. We will discuss it more thoroughly in Section VII.

### v5

@@ III-C Existence of the Codebooks and Achievability Bound
    $\begin{split}\log M_{n}=&nC(\mu,\Psi(n))+n\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))\\ +&\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}\\ \times&\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\\ +&O(\log n).\end{split}$ (110)
@@
    $\gamma_{1}=n\mathsf{C}(\mu,\Psi,\sigma^{2})-n(1+\rho)\hat{\mathsf{C}}(\Psi(n),\mu\Psi(n),\sigma^{2})$ (112)
@@
    The power level $\mu\Psi(n)=c_{4}\cdot\frac{\sqrt{n\delta}}{n}$ with $c_{4}$ is any constant such that
    $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})<\delta.$ (113)
Under the above conditions, with sufficiently large $n$, we can choose $c_{0}=13$, $c_{1}=1$, $c_{2}=11$, $c_{3}=4$, and $c_{5}=3$ and a proper power level $c_{4}=\frac{2\sigma^{2}}{\mu}\sqrt{\delta\ln 2}$. We obtain
$\log M^{*}(n,\epsilon,\sigma,\delta)$ $\geq\mu\sigma^{2}(\delta n\log e)^{1/2}$
$+\sigma(2\mu)^{1/2}(n\delta)^{1/4}$ $(\log e)^{3/4}\Phi^{-1}(\epsilon)+o(n^{\frac{1}{4}}),$ (114)
@@
$\log K(n,\epsilon,\sigma,\delta)=\begin{cases}\frac{1-\mu^{2}\sigma^{2}}{\mu}(\delta n\log e)^{1/2}+O(n^{1/4})\\ \,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<\frac{1}{\mu^{2}},\\ O(1)\,\ \ \ \ \ \,\text{otherwise}.\end{cases}$ (115)
@@
For each index $s$, we expurgate the $(1-\lambda_{2})M_{n}$ codewords with largest $\varepsilon_{sw}^{(2)}$ for from each sub-codebook $\{\bm{x}_{sw},w\in\langle M\rangle\}$. Thus,we get a new family of codebooks, each of which has size $\lambda_{2}M$. Without loss of generality, we assume that the remaining codewords have indexes from $1$ to $\lambda_{2}M$, and denote the maximum of error probability of the second type in these new codes as $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. These codewords with indexes from $\lambda_{2}M+1$ to $M$ have error probabilities of the second type at least larger than $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. As the average error probability of the second type for each old codebook is less than $\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. We have for each $s$, the following inequality holds,
$\begin{split}&\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)\cdot M_{n}\\ \geq&\sum_{w=1}^{M}\varepsilon_{sw}^{(2)}=\sum_{w=1}^{\lambda_{2}M}\varepsilon_{sw}^{(2)}+\sum_{w=\lambda_{2}M+1}^{M}\varepsilon_{sw}^{(2)}\\ \geq&(1-\lambda_{2})M_{n}\cdot\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}.\end{split}$ (116)
Thus, the new family of codebooks should satisfy
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(2)}\leq\frac{\lambda_{1}}{1-\lambda_{2}}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg).$ (117)
For the condition of reliability, for sufficiently large $n$, the following inequality on the maximal error probability of the code family holds,
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(2)}\leq$ $\frac{1}{n^{c_{0}-c_{1}-c_{2}}}\cdot\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$
$\overset{(a)}{\leq}$ $\frac{2}{n^{c_{0}-c_{1}-c_{2}}}$
$\leq$ $\epsilon-\epsilon_{1}$ (118)
where the step (a) follows from Lemma 3. Thus, we get the maximal probability of error satisfies
$\underset{s,w}{\max}\,\ \,\varepsilon_{sw}^{(2)}+\varepsilon_{sw}^{(1)}\leq\epsilon-\epsilon_{1}+\epsilon_{1}<\epsilon.$ (119)
For the covertness condition in (108), from (113) and Lemma 3, the term $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ satisfies
$(1+\iota^{\prime})$ $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
@@
$\log$ $\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big)+An\cdot ae^{-bn}$
$+$ $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})+\lambda_{3}+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
@@
$+$ $\frac{1}{n^{c_{3}}}+(1+\iota^{\prime})\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$\leq$ $\delta.$ (121)
@@
$\exp$ $\bigg[-M\cdot\bigg(\frac{\lambda_{2}\lambda_{3}^{2}\sigma^{4}}{8n^{2}\Psi^{2}(n)}-\mathbb{H}_{b}(\lambda_{2})\bigg)\bigg]+\frac{1}{\lambda_{1}}$
@@
$\leq$ $1.$ (126)
$\log M^{*}(n,\epsilon,\delta)\geq$ $\frac{n}{2}\log(1+\mu c_{4}\frac{\sqrt{n}}{n})+\frac{(\mu^{2}-\mu)c_{4}\frac{\sqrt{n}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})}\cdot\log e+\sqrt{n\cdot\frac{2\mu^{2}c_{4}\frac{\sqrt{n}}{n}+\frac{c_{4}^{2}\mu^{2}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})^{2}}}\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\log e$
@@
With $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get a lower bound of $\log M^{*}(n,\epsilon,\sigma,\delta)$ in (III-C). The condition in (111) is equivalent to
$\log M+\log K=\gamma_{1}+c_{5}\log n+1$ (128)
Consequently, the least number of bits of $K$ should satisfy
$\log K$
$=$ $\max\{0,\gamma_{1}+c_{5}\log n+1-\log M\}$
$\overset{(a)}{=}$ $\max\{0,n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))$
- $\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]$
- $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}$
@@
$\overset{(b)}{=}$ $\max\{0,n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))$
- $\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})-\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]$
@@
$\begin{split}&n[\mathsf{C}(\mu,\Psi(n),\sigma^{2})-C(\mu,\Psi(n))]\\ =&n[\frac{1}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\frac{1}{2}\log(1+\mu\Psi(n))]\\ =&\frac{n}{2}\log\big[1+\frac{\mu\Psi(n)(1-\sigma^{2})}{\sigma^{2}(1+\mu\Psi(n))}\big]\\ =&\frac{1}{2}\mu\big(\frac{1}{\sigma^{2}}-1\big)c_{4}\sqrt{n}\log e+O(1).\end{split}$ (130)
In addition,
$\begin{split}&n[\hat{\mathsf{C}}(\Psi,\mu\Psi(n),\sigma^{2})+\bar{\mathsf{C}}(\mu^{2}\Psi(n),\mu\Psi(n))]\\ =&\frac{(\mu-1)\Psi(n)(1+\mu\Psi(n)+\mu\sigma^{2}+\mu^{2}\Psi(n))}{2(1+\mu\Psi(n))(\sigma^{2}+\mu\Psi(n))}\log e\end{split}$ (131)
By using $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$ in (130) and (131), the key size in (129) is rewritten as
$\max\{0,\frac{1}{2}\mu\big(\frac{1}{\sigma^{2}}-1\big)c_{4}\sqrt{n}\log e$
- $\frac{n(\mu-1)(1+\mu\sigma^{2})c_{4}\sqrt{n}}{2\sigma^{2}}\}$
$=$ $\max\{0,\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}c_{4}\sqrt{n}\log e+O(1)\}$ (132)
and the term $\sqrt{n\mathsf{V}(\Psi(n),\mu\Psi(n))}=O(n^{1/4})$. Let $\phi(\mu,\sigma)=\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}$, we discuss it in different cases.
  • If $\sigma^{2}\leq\frac{1}{\mu^{2}}$, i.e., the channel between Alice and Willie is less noisy than the main channel, we have
    $\log K=\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}c_{4}\sqrt{n}\log e+O(n^{1/4}).$ (133)
  • If $\sigma^{2}>\frac{1}{\mu^{2}}$, then $\phi(\mu,\sigma^{2})<0$ and the first order asymptotics of the optimal number of key bits should be $O(1)$.
@@
$\log K=\begin{cases}\frac{1-\mu^{2}\sigma^{2}}{2\sigma^{2}}\cdot c_{4}\sqrt{n}\log e+O(n^{1/4}),\\ \,\ \ \ \ \ \ \ \ \ \ \ \,\text{if}\,\ \,\sigma^{2}<\frac{1}{\mu^{2}}\\ O(1)\,\ \ \ \ \ \,\text{otherwise}\end{cases}$ (134)
Finally, we consider the inequality (113)
$\delta$ $>\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$
$=\frac{n}{2}\left[\frac{\mu\Psi(n)}{\sigma^{2}}-\ln(1+\frac{\mu\Psi(n)}{\sigma^{2}})\right]$
$\geq\frac{1}{4}n\frac{\mu^{2}\Psi^{2}(n)}{\sigma^{2}}$ (135)
and we get a proper $c_{4}$ as
$c_{4}=\frac{2\sigma^{2}}{\mu}\sqrt{\delta\ln 2}.$ (136)
Combing (III-C), (134) and (136), we finally get the size of the $M$ in (6) and the size of the key in (115). ∎
@@
The support set of the codewords is $\{\bm{x}:\sqrt{\mu^{2}n\Psi(n)}\leq\|\bm{x}\|\leq\sqrt{n\Psi(n)}\}$. The upper bound of $\|\bm{x}\|$ is for the power constraint, and we set the inner bound $\sqrt{\mu^{2}n\Psi(n)}$ here for lower bounding the decoding error probability. If there exists codewords with too small $l^{2}$ norm, then there exists no $\gamma$ to lower bound the first type of error in decoding. In fact, the bounding of the first type of error in Lemma 4 relies on the fact that the power of each codeword lies in a finite interval (See Remark 5). Here we choose truncated Gaussian with decreasing variance $\mu\Psi(n)$. The parameter $\mu$ acts as both the parameter of variance and the parameter of truncation, which makes the analysis more convenient and the analytic formulas more concise. The truncation parameters for the inner bound and outer bound of $l^{2}$ norm of the codewords can be different, which will lead to the first and second order asymptotics with different coefficients. Nevertheless, it will not change the order of them as $O(n^{\frac{1}{2}})$ and $O(n^{\frac{1}{4}})$. In addition, the truncation parameter of the outer bound of radius will be reflected in the coefficient of these asymptotics as the power constraint is necessary in AWGN channels.
@@
Note that the truncation parameter $\mu$ lies in both the size of codebooks (6) and the size of the key (115), which is in contrast with the analysis of the first order asymptotics of covert communication over AWGN channels [10] [11] [12]. If we apply random coding with Gaussian distribution as in the achievability part of Theorem 5 in [12] or Theorem 9.1.1 in [38], we can not guarantee that the distribution of the codebook is Gaussian since we can not use $n\rightarrow\infty$ in the finite blocklength regime. The advantages of truncated Gaussian family as the generating distribution of the codewords are as follows. Firstly, all codewords satisfying power constraint is ensured so that we can control the power level easily. Secondly, the truncated Gaussian family is close to Gaussian family, and the effect of truncation on the output distribution is convenient to manipulate to satisfy the covert constraint.
Remark 14.
To satisfy the covert constraint with KL divergence in (12), the power level should be $\Psi(n)=O(\frac{1}{\sqrt{n}})$ in (113) of the condition (d). The kernel lies in the divergence between two Gaussian distributions $Q^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},0}$. If the covert constraint is in the form of other divergences with proper upper bound $\delta$, such as total variational distance, Hellinger distance or $\chi^{2}$-divergence, the results will be almost the same except that $\Psi(n)$ may have a different constant coefficient. This fact will be illustrated later in Section V.
Remark 15.
For covert communication over AWGN channels, previous works [10] [11] have established the characterization of the first order asymptotics of the throughput as $O(\sqrt{n})$. This work focuses on refining these results by characterizing both the first and second order asymptotics as functions of the blocklength $n$, decoding error probability $\epsilon$, the background noise level parameter $\sigma$ and covert parameter $\delta$, which is not reported in [10] [11]. The analytic investigation stresses the influence of the ratio of the noise level (which is $\sigma^{2}/1$ in this work) on the throughput and the size of key bits.
Remark 16.